A research worker wishes to estimate the mean of a population by using sufficiently large sample. The probability is 0.95 that the sample mean will not differ from the true mean by more that 25% of the standard deviation. How large sample should be taken?


1
Expert's answer
2021-03-17T15:34:57-0400

Solution:

We have "P(|\bar{X}-\mu|<\sigma / 4)=0.95"

Divide by "\sigma / \sqrt{n}"

"{P}\left(\frac{\bar{X}-\mu}{\sigma / \sqrt{n}}<\frac{\sqrt n}{4}\right)=0.95"

But "\frac{\bar{X}-\mu}{\sigma / \sqrt{n}}" has a standard normal distribution.

And, "{P}(-1.96< \text{Standard Normal} <1.96)=0.95"

So, "1.96=\sqrt{n} / 4"

"\Rightarrow 1.96\times4=\sqrt n
\\ \Rightarrow n=(7.84)^2=61.4656
\\\Rightarrow n=62"

So, a sample of 62 should be taken.